for what value of a would the following system of equations have an infinite number of solutions?
2x - y = 8
6x - 3y = 41
A: 2
B: 6
C:8
D:24
E:32

Answer by robertb(5830)

Where's the a attached to?
Gonna get back to you later... Got a class to teach...
I'm back...
2x - y = 8
6x - 3y = 4a.
Multiply the top equation by 3, to give 6x - 3y = 24. This is almost identical with the bottom equation, except for constants. Forthe system to have an infinite number of solution, the constants must be equal. Hence, 4a = 24, and thus a = 6. Have a good one...